Use the spinner below to answer the question. Write the probabilities in percent notation rounded to the nearest tenth. Assume that it is equally probable that the pointer will land on any one of the eight numbered spaces. If the pointer lands on a borderline, spin again.

WON'T LET ME ADD FOR SOME REASON. SPINNER IS NUMBERED 1-8. 1,7,5,3 YELLOW AND 2,8,6,4 IS BLUE

Write the probabilities in percent notation rounded to the nearest tenth.

(a) The probability that the arrow will land on even number is Answer

%

(b) The probability that the arrow will land on odd number is Answer

%

(c) The probability that the arrow will land on blue field is Answer

%

(d) The probability that the arrow will land on odd number and blue field is Answer

%

(e) The probability that the arrow will land on odd number or blue field is Answer

%

(f) The probability that the arrow will land on a number greater than 6 is Answer

%

(g) The probability that the arrow will not land on a number smaller than 3 is Answer

%

(i) The probability that the arrow will land on 5 or even number is Answer

%

(j) The probability that the arrow will land on 5 or prime number is Answer

%
Answer #1

Probability is the numerical value given to the chance of an event happening.

Probability = No. of desired outcomes/ Total outcomes

Here total outcomes = 8

Here we have  1,7,5,3 YELLOW AND 2,8,6,4 IS BLUE

That means there 4 odds and 4 even

Colour wise there are 4 yellow and 4 blues

(a) The probability that the arrow will land on even number :

P(even ) = No. of even / 8

=4 / 8

Ans: 0.5 = 50%

(b) The probability that the arrow will land on odd number is

P(odd ) = No. of odd / 8

=4 / 8

Ans: 0.5 = 50%

(c) The probability that the arrow will land on blue field :

P(blue) = No. of blue / 8

=4 / 8

Ans: 0.5 = 50%

(d) The probability that the arrow will land on odd number and blue field

Blue and odd numbers are :None

Since all odds are yellow

P(blue and odd) = 0 / 8

ans: 0

(e) The probability that the arrow will land on odd number or blue field is

P(Odd or blue) = P(odd) + P(blue) - P(blue and odd) .........addition theorem of probability

=0 .5 + 0.5 - 0

ans: 1

(f) The probability that the arrow will land on a number greater than 6 is

P(greater than 6) are either 7 or 8 so 2 option

P(greater than 6) = 2 / 8

Ans: 0.25 = 25%

(g) The probability that the arrow will not land on a number smaller than 3

not smaller than 3 means 3 or greater than 3

3 or greater than 3 : 3 ,4,5,6,7,8 =6 options

P(not land on a number smaller than 3) = 6 / 8

ns: 0.75 = 75%

(i) The probability that the arrow will land on 5 or even number

P(5 or even) = P(5) + P(even) + P(5 and even)

= 1 / 8 + 0.5 -0

Ans: 0.625 = 62.5%

5 is single number so only one option. For 5 and even, they both are mutually exclusive that is can't happen together so p = 0.

(j) The probability that the arrow will land on 5 or prime number

Prime numbers are : 2,3,5,7 : 4 options

5 and prime : common between 5 and prime: 5: 1 option

P(5 or prime) = P(5) + P(prime) - P(5 and prime)

=1 / 8 + 0.5 - 1 /8

ans: 0.5 = 50%